OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} ((x+1)(x+2)(x+3)(x+4))^{1/4}-x\]

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

@hartnn ?

OpenStudy (anonymous):

@DLS ? @Azureilai ? @bcr1997 ? @chmvijay ? @Hero ? @insa ? @karlacampos ? @matricked ?

OpenStudy (anonymous):

should i tell my approach?

OpenStudy (akashdeepdeb):

Go ahead.

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}x[(1+1/x)+(1+2/x)+(1+3/x)(1+4/x)]^{1/4}-x\]

OpenStudy (anonymous):

sorry i think i wrote wrong instead of + sign there should be multiplication

OpenStudy (anonymous):

\lim_{x \rightarrow \infty}x[(1+1/x)(1+2/x)(1+3/x)(1+4/x)]^{1/4}-x

OpenStudy (akashdeepdeb):

Perfect! Now, \[\lim_{x \rightarrow \infty} {(((x+1)(x+2)(x+3)(x+4))}^{\frac{1}{4}} - x)\] Now after you take x common from all the terms, you have \((x^4)^{\frac{1}{4}} = x\) Then you'd get: \[\lim_{x \rightarrow \infty} x((1 + \frac{1}{x})(1 + \frac{2}{x})(1 + \frac{3}{x})(1 + \frac{4}{x}) - 1)\] Use the property of limits: \[\lim_{x \rightarrow a} (f(x) ~~.~ ~g(x)) = \lim_{x \rightarrow a} f(x) ~~. ~~\lim_{x \rightarrow a} g(x)\] You'll get: \[\lim_{x \rightarrow \infty} x ~~. ~~\lim_{x \rightarrow \infty} ((1+ \frac{1}{x})(1 + \frac{2}{x})(1 + \frac{3}{x})(1 + \frac{4}{x}) - 1)\] Which gives you: \(\infty . 0\) = 0 ? Something does not seem right or is it? @hartnn @ganeshie8

OpenStudy (akashdeepdeb):

I missed out the \(\frac{1}{4}\) above. Even if it was there, it wouldn't have made any difference I suppose.

OpenStudy (akashdeepdeb):

I am very sure, 0 is not the answer. Hold on let me try again.

OpenStudy (akashdeepdeb):

Do we have to use binomial Theorem by any chance? Yeah, even I checked on wolfram, but couldn't check the step-by-step solution.

OpenStudy (ikram002p):

got a headach trying to expand the LH @_@

OpenStudy (akashdeepdeb):

Got the answer? O.O

ganeshie8 (ganeshie8):

\[\lim_{x \to \infty} ~\left((x+1)(x+2)(x+3)(x+4)\right)^{1/4}-x\] \[\lim_{t \to 0} ~ \dfrac{\left((1+t)(1+2t)(1+3t)(1+4t)\right)^{1/4}-1}{t}\] indeterminate form, apply L'hosps \[\lim_{t \to 0} ~ \dfrac{1}{4}\left((1+t)(1+2t)(1+3t)(1+4t)\right)^{-3/4}\times \left(1 + 10t + \mathcal{O(t^2)} \right)'\] \[\lim_{t \to 0} ~ \dfrac{1}{4}\left((1+t)(1+2t)(1+3t)(1+4t)\right)^{-3/4}\times \left(0 + 10 + \mathcal{O(t)} \right)\] \[ \dfrac{1}{4}\left((1+0)(1+0)(1+0)(1+0)\right)^{-3/4}\times \left(0 + 10 + 0 \right)\] \[\dfrac{5}{2}\]

OpenStudy (akashdeepdeb):

Awesome! :'O

OpenStudy (akashdeepdeb):

hahahahahahah! :D

OpenStudy (ikram002p):

me :- one step late *duh*

ganeshie8 (ganeshie8):

lol that suggestion of t->0 was partly given by @DLS

OpenStudy (ikram002p):

* going to kill myself * lolz , it really cuz a headach xD

OpenStudy (ikram002p):

need to review calcules ,no doubt thats it xD

OpenStudy (akashdeepdeb):

I wanted to use, L'Hospital. :'|

OpenStudy (dls):

<3 <3

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

how did u transform it to the 2nd line

OpenStudy (anonymous):

@ganeshie8 ?

ganeshie8 (ganeshie8):

substitute 1/x = t

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} ((1+\frac{ 1 }{ t })(1+2/t)(1+3/t)(1+4/t))^{1/4}-1/t\]

ganeshie8 (ganeshie8):

added few more lines : \[\lim_{x \to \infty} ~\left((x+1)(x+2)(x+3)(x+4)\right)^{1/4}-x\] \[\lim_{x \to \infty} ~x \left[\left((1+\frac{1}{x})(1+\frac{2}{x})(1+\frac{3}{x})(1+\frac{4}{x})\right)^{1/4}-1\right]\] sub \(\frac{1}{x} = t\) as \(x \to \infty , t \to 0\) \[\lim_{t \to 0} ~ \dfrac{\left((1+t)(1+2t)(1+3t)(1+4t)\right)^{1/4}-1}{t}\] indeterminate form, apply L'hosps \[\lim_{t \to 0} ~ \dfrac{1}{4}\left((1+t)(1+2t)(1+3t)(1+4t)\right)^{-3/4}\times \left(1 + 10t + \mathcal{O(t^2)} \right)'\] \[\lim_{t \to 0} ~ \dfrac{1}{4}\left((1+t)(1+2t)(1+3t)(1+4t)\right)^{-3/4}\times \left(0 + 10 + \mathcal{O(t)} \right)\] \[ \dfrac{1}{4}\left((1+0)(1+0)(1+0)(1+0)\right)^{-3/4}\times \left(0 + 10 + 0 \right)\] \[\dfrac{5}{2}\]

OpenStudy (anonymous):

what is this (1+10t ....(t^2) ??

ganeshie8 (ganeshie8):

expand the product \((1+t)(1+2t)(1+3t)(1+4t)\)

ganeshie8 (ganeshie8):

Clearly, the constant term is 1x1x1x1 = 1

ganeshie8 (ganeshie8):

the coefficient of t = (1 + 2 + 3 + 4) = 10

ganeshie8 (ganeshie8):

we don't need higher order terms cuz they vanish in the next line when u take the limit.

ganeshie8 (ganeshie8):

if thats confusing, you may simply expand the whole thing and take the derivative ^

OpenStudy (anonymous):

thanks i got it!

OpenStudy (anonymous):

:)